### v1

# Making Language Model a Hierarchical Classifier and Generator

Abstract
Decoder-only language models, such as GPT and LLaMA, generally decode on the last layer. Motivated by human’s hierarchical thinking capability, we propose that a hierarchical decoder architecture could be built with different layers decoding texts simultaneously. Due to limited time and computationally resources, we choose to adapt a pretrained language model into this form of hierarchical decoder. Language heads of the last layer are copied to different selected intermediate layers, and fine-tuned with different task inputs. By thorough experiments, we validate that these selective intermediate layers could be adapted to speak meaningful and reasonable contents, and this paradigm of hierarchical decoder can obtain state-of-the-art performances on multiple tasks such as hierarchical text classification, classification-guided generation, and hierarchical text generation. HdLM outperforms all baselines on WoS, DBpedia, ESconv, EmpatheticDialogues, and several cognitive tests. We also provide thorough theoretical analysis to validate the convergence and computational savings of our methodology. This study suggests the possibility of a generalized hierarchical reasoner, pretraining from scratch.
1 Introduction
Modern Large language models (LLM), such as GPT [1] and Llama [2], have made remarkable progresses on natural language tasks [3]. Based on the decoder-only architecture, such language models exhibit impressive generalizability and scalability on different tasks and scenarios. Among these tasks, textual reasoning is always challenged which often requires the model to explicitly plan the immediate steps and bootstrap the long-term rewards. To enhance such capabilities, chain-of-thought (CoT) [4] and corresponding finetuned or test-time scaling methods [5; 6; 7] are proposed, with the thought formulated either on the token-space or the latent-space.
However, these reasoning LLMs do not have an explicit hierarchical thinking mechanism, therefore might suffer from suboptimal solutions on complex reasoning tasks [8]. Their reasoning capabilities are constrained from the scale of annotated data, and the test-time scaling might be upper-bounded [9]. On the other hand, mankind is naturally empowered with hierarchical thinking capabilities [8], typically in two aspects:
(i) Different abstraction of conceptual levels: the coarse-grained, strategic decisions guide the generation of fine-grained, detailed decisions.
(ii) Sequential determination: a key determination is first selected from available options while subsequent options are subject to the selection.
Motivated from these hierarchical thinking mechanisms, models with automatic hierarchical thinking should be explored. Similar studies were mainly based on encoder-decoder architectures and may be classified into model-wise, scale-wise and layer-wise prototypes [10].
Figure 1: Paradigm of HdLM compared to a conventional decoder-only language model. In contrast to the last layer decoding, HdLM employs different layers to decode different levels of texts.
In this paper, we propose a new type of decoder called Hierarchical decoding Language model (HdLM), combining the model-wise and layer-wise prototypes. We inherit the strong semantic understanding and generation capabilities of modern LLM, but allow its multiple layers to decode plausible textual contents. An earlier layer may produce more prerequisite, coarse-grained expressions which also work for the context of resulted, fine-grained expressions, which are decoded by later layers (Figure 1). Such type a language model can automatically produce reasonable performance on sequential pipelines of classification and generation tasks, which are verified by our substantial experiments. Comparing with the conventional reasoning LLM, HdLM also enjoys computational savings for both training and inference, which are both verified by theoretical derivation and empirical observations. Our code and model can be found on https://github.com/coswindywang/HdLM. Our main contributions can be summarized as follows:
(1) We propose HdLM, which can deal with different types of hierarchical tasks, including hierarchical text classification, classification-guided generation, and hierarchical text generation.
(2) We provide a theoretical analysis on the computational benefit HdLM, and also a discussion on its convergence.
(3) We conduct substantial experiments and corresponding analysis to verify HdLM’s performance, on both classification and generation metrics, as well as its out-of-domain generalization capability.
(4) We discuss about its connection with alternative solutions such as vanilla SFT, cascade, multi-hop, and its scalability and the possibility of an end-to-end framework.
2 Method
In this section, we first formalize the problem, then propose a dual-layer fine-tuning mechanism, and finally a two-pass inference paradigm.
2.1 Problem Formulation
Given a language model $\mathcal{M}$, its total number of layers is $K$, and the language head is $\mathcal{H}$. Different from traditional query-response tasks, here we try to define a generalized hierarchical textual tasks. Given a user query $q$, hierarchical textual tasks requires the agent to generate a sequence of responses $\mathbf{r}_{1:D}:=\{r_{1},r_{2},\cdots,r_{D}\}$, in which $D$ is the hierarchical depth. The standard format of hierarchical textual data then becomes $(q,\mathbf{r}_{1:D})$. We use $L$ and $L_{d}$ to denote the lengths of $q$ and $r_{d}$, respectively.
The above hierarchical textual task can be solved recursively. That is, for each step $i\in\{1,2,\cdots,D\}$, the current response can be produced grounded by the query and prior responses:
$\mathcal{T}_{i}\in\{\mathcal{C},\mathcal{G}\}:r_{i}\leftarrow\mathcal{M}(q,\mathbf{r}_{1:i-1})$ (1)
where $\mathcal{T}$ denotes a uni-step subtask. Here we further argue $\mathcal{T}$ can be generally classified into two categories: the classification task $\mathcal{C}$ and the generation task $\mathcal{G}$. As indicated in Figure 2, the following different paradigms can be summarized from hierarchical textual tasks:
Hierarchical text classification (HTC):
classifications are asked from the coarse-grained label to the fine-grained label. The lower-level label candidates are usually constrained by the choice of the higher-level label. We denote this paradigm as $\mathcal{C}\rightarrow\mathcal{C}$.
Classification-guided generation (CgG):
a classification result is first required, then the final response is generated based on that classified label. This paradigm can be denoted as $\mathcal{C}\rightarrow\mathcal{G}$.
Hierarchical text generation (HTG):
the final response is posterior inferenced by intermediate generations, such as the chain of thought, reasoning, or thoughts. This paradigm is denoted as $\mathcal{G}\rightarrow\mathcal{G}$.
Other hierarchical tasks:
there might be even more completed hierarchical tasks. For example, the generative classification ($\mathcal{G}\rightarrow\mathcal{C}$) and ‘think, classify and act’ ($\mathcal{G}\rightarrow\mathcal{C}\rightarrow\mathcal{G}$). We leave these types of tasks for future works.
While there are evident methods to solve hierarchical textual tasks, either by computation time scaling (such as CoT or reasoning LLM), or the multi-hop inference, in this paper we propose a single-model framework which has less computational overheads.
Figure 2: Typical paradigms (hierarchical text classification, classification-guided generation, hierarchical text generation) of hierarchical text tasks. $\mathcal{C}$ denotes classification and $\mathcal{G}$ denotes generation.
2.2 Architecture
In this section, we propose an adjusted architecture based on the decoder-only transformer, which can also deal with hierarchical textual tasks by post-hoc adaptations. We first make a reasonable assumption that the depth of hierarchical textual task is smaller, i.e., $D<K$. Then we select $D-1$ intermediate layers with their indices satisfying:
$k\in\{k_{1},\cdots,k_{D-1},K\},0<k\leq K$ (2)
These layers, along with the final layer, are used to decode $D$ responses. To achieve this objective, we replicate the language heads in the $K$-th layer to the $D-1$ intermediate layers, with the parameter randomly initialized.
$\mathcal{H}_{d}\leftarrow\mathcal{H}_{K},d=1,2,\cdots,D-1$ (3)
To make the derivation clearer, we propose some notations. Given a language model $\mathcal{M}$, $\mathcal{M}_{k}$ represents its $k$-th layer. The forward pass from the $k_{1}$ to $k_{2}$-th layer is denoted by $\mathcal{M}_{k_{1}:k_{2}}$, while $\mathcal{M}_{k_{1}:k_{2}}^{\xrightarrow{L_{1}}}$ denotes the forward pass plus a textual generation.
The final layer still holds the standard supervised Fine-Tuning (FT) loss:
$\mathcal{L}^{\text{FT}}_{D}=-\frac{1}{L_{D}}\sum_{j=1}^{L_{D}}\log\left[\text{P}(r_{D}(j)|q,\mathbf{r_{1:D-1}},r_{D}(1:j-1))\right]$ (4)
where $j$ denotes the $j$-th token of $r_{D}$.
2.3 Training
The $D-1$ intermediate layers can not decode reasonably with newly added heads. To educate them to decode proficient language, post-hoc adaptation is needed.
Given the hierarchical textual sample $(q,r_{0:d})$, the latent vector of each hierarchical level can be calculated recursively by the forward pass of each LM block:
$e_{1}$ $=\mathcal{M}_{0:k_{1}}(q),\quad e_{d+1}=\mathcal{M}_{k_{d}:k_{d+1}}(e_{d},r_{d}),\quad\mathcal{L}_{d}=\mathcal{L}^{\text{FT}}(q,e_{1:d},r_{1:d}),\quad d=1,2,\cdots,D$ (5)
with the finetuning loss $\mathcal{L}_{d}$ implemented on each decoding layer. The final loss can then be the linear combination of them: $\mathcal{L}=\sum_{d=1}^{D-1}f_{d}\mathcal{L}_{d}+\mathcal{L}_{D}$ with the list of $\{f_{1},\cdots,f_{D-1},1\}$ as the loss weights. During the training, we implement the causal mask for both prior tokens and subsequent responses. Figure 3 exhibits the detailed masking matrices for all decoding layers.
Figure 3: The pipeline of HdLM, with exampled $D=3$ and $K=5$. (a): A depth=3 hierarchical task on a conventional LM. (b): Replicate the language heads of the $K$-th layer to layer $k_{1},k_{2}$. (c): Training HdLM with the fine-tuning loss of $r_{d}$ on the $k_{d}$-th layer. (d): Inference HdLM on the $k_{d}$-th layer sequentially with decoded $\hat{r}_{d}$ and corresponding length $\hat{L}_{d}$.
2.4 Inference
We keep a similar forward logic of inference to the training. For the query $q$, we still calculate its logits throughout all attention layers, similar with the conventional language model. For each $k$-th layer, the prior latent vector is employed to decode the $k$-th response, while the new decoded latent is also employed to forward pass the final layer, until the final response is decoded grounded by all previous latents.
$e_{1},\hat{L}_{1}$ $=\mathcal{M}_{0:k_{1}}^{\xrightarrow{L_{1}}}(q),\quad e_{d+1},\hat{L}_{d+1}=\mathcal{M}_{k_{d}:k_{d+1}}^{\xrightarrow{L_{d}}}(e_{d})$ (6)
$\hat{r}_{d}$ $=\mathcal{H}_{k_{d}}(e_{d}(L+\sum{\hat{L}_{<d-1}}:)),\quad d=1,2,\cdots,D$ (7)
Note the above expression has a similar structure to the state-space model (SSM).
3 Theoretical analysis
This section theoretically analyzes the computation savings of HdLM compared to a conventional LLM. We then briefly discuss the training convergence of HdLM with the connection between the index of the first decoding layer and the task depth.
3.1 Computational complexity analysis
In this subsection, we compare the computation FLOPs of our HdLM with a standard LLM, and verify that HdLM results in both training and inference reduction. To make the formulation simple and clear, here we discuss the 2-depth hierarchical task, i.e., $D=2$.
Parameter definitions.
The input sequence length $L$, the output sequence lengths $L_{1}$ and $L_{2}$, the hidden dimension $E$, indexes of two decoding layers $k_{1}<K$ and $k_{2}=K$, and the expand ratio of FFN dimension is $c$. We use $\mathcal{F}$ to denote the FLOPs.
Training FLOPs savings.
We begin with the forward FLOPs of a standard transformer from Equation 20 derived in Appendix: $\mathcal{F}_{\text{train}}\rightarrow 3(8+4c)E^{2}KL$  [^1] Here we omit the batch size term $B$ since different samples’ computation within a batch is parallel.. The LLM baseline simply treats the input, level-1 output and level-2 output as a flat sequence, with all layers engaged:
$\mathcal{F}_{\text{train}}^{\text{baseline}}\approx 3fK(L+L_{1}+L_{2})$ (8)
where we use $f$ to represent the fixed coefficient $(8+4c)E^{2}$ for simplicity. One the other hand, the training FLOPs of HdLM can be expressed as
$\mathcal{F}_{\text{train}}^{\text{{HdLM}}}\approx 3fk_{1}(L+L_{1})+3f(K-k_{1})(L+L_{1}+L_{2})$ (9)
The training FLOPs savings then becomes
$\mathcal{F}_{\text{train}}^{\text{baseline}}-\mathcal{F}_{\text{train}}^{\text{{HdLM}}}\approx 3fk_{1}L_{2}>0$ (10)
Inference FLOPs savings.
The inference FLops of LLM baseline can be expressed as
$\mathcal{F}_{\text{infer}}^{\text{baseline}}=fK\sum_{j=1}^{L_{1}+L_{2}}(L+j-1)$ (11)
from Equation 21 derived in Appendix. While for HdLM, the inference FLOPs becomes
$\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}=fk_{1}\sum_{j=1}^{L_{1}}(L+j-1)+f(K-k_{1})\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)$ (12)
and it is easy to further prove $\mathcal{F}_{\text{infer}}^{\text{baseline}}-\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}>0$ (detailed derivation in Appendix).
3.2 Convergence of decoder with arbitrary layers
This subsection briefly discusses the convergence issue, i.e., how can we ensure any implementation decoding layer (with the index of $k_{d}$), has enough depth to learn the target semantic hierarchy (with depth $d$)? Our derivation is mainly motivated by the Minimum number of layers Theorem proposed in [11]
Theorem 1.
The last position in the sequence cannot retrieve the embedding vector $x_{D}$ of the target token with $K$ transformer layers if $K<\log_{3}(2D)$.
Due to the page limit, here we leave the detailed assumption in [11] and the connection between retrieval and classification in the Appendix. Based on its conclusion, here we prove our Corollary 2:
Corollary 2.
Assume Theorem 1 holds in $\mathcal{C}$ or $\mathcal{G}$, then for any $d\leq D$, the first $k_{d}$ layers of transformer can always learn the top $d$-depth task information.
Proof.
Given the strict increasing, non-repeat integer sequence $\{k_{d}\}$ with $d\in\{1,2,\cdots,D\},k\in\{1,2,\cdots,K\},D\leq K$, it is obvious to have $k_{d}\geq d$. Since $3^{d}>2d$ holds for $d>=1$, then we have $k_{d}\geq d>\log_{3}{2d}$, which contradicts the situation in Theorem 1. ∎
Corollary 2 ensures the theoretical feasibility of our approach. In the next section, we also verify this conclusion from empirical results.
4 Experiment
In this section, we consider answering the following research questions: RQ1: If the training of HdLM can converge with an arbitrary setting of $k$? RQ2: Can HdLM perform well on different types of hierarchical textual tasks and outperform conventional baselines? RQ3: Can HdLM generalized well on out-of-domain hierarchical textual datasets? RQ4: How to determine the key parameters and different components of HdLM? RQ5: What is the connection of HdLM with previous standardized or test-time scaling baselines, such as reasoning, multi-hop, or cascade solutions?
To address these issues, we first list the datasets we use for different types of hierarchical tasks, then the experimental settings and evalulation metrics, and followed by the main results. We finally provide thorough discussion on the open-domain performance, the ablation and sensitivity studies, and the actual computational savings.
4.1 Datasets
HTC ($\mathcal{C}\rightarrow\mathcal{C}$).
We use the famous WoS [12] (with depth $D=2$) and DBpedia [13] (with depth $D=3$) as the training datasets. WoS are abstracts of published papers from Web of Science while DBpedia extracts structured information from Wikipedia.
CgG ($\mathcal{C}\rightarrow\mathcal{G}$).
We employ the emotional support conversation (ESC) scenario, in which each turn of conversation is annotated with the user’s emotion and the assistant’s response strategy. The model needs to sequentially classify the user emotion or the response strategy to further enhance the response quality. We utilize ESconv [14] as both training and in-domain (ID) test sets. To further validate the response generation, EmpatheticDialogues [15] is employed as an out-of-domain (OOD) test set which lacks of the strategy annotation.
HTG ($\mathcal{G}\rightarrow\mathcal{G}$).
Here we examine two typical famous theory-of-mind (ToM) tests, the Sally-Anne false belief experiment, and the Vignette-based Problem. These cognitive tests contain misleading or disturbing information and the agent needs to reason from the ego-centric perspective. In these scenarios, the response is the (thought, answer) pair with depth $D=2$. We train HdLM on the ToMI [16] and BigToM [17] benchmarks, while zero-shot evaluate the model on the Vignette-based test provided by [18].
4.2 Setting
Implementation Details.
We conduct a post-hoc adaptation on the basis of Llama3-8B-Instruct [19], which has a total $K=32$ attention layers. During training, we use the AdamW optimizer with decay of 0.01 and the cosine scheduler. The training batch size is 16 and the sequence length is 2048. The experiment is running on LlamaFactory [20] with 32 A100 GPUs, lasting about 16 hours. Other settings, which is dataset-wise, are listed in the Appendix.
Classification metrics.
F1-related scores including Micro-F1 and Macro-F1 are used. Micro-F1 considers the overall precision and recall of all instances, while Macro-F1 equals the average F1-score of labels. For CgG tasks, we also provide the classification accuracy, and the preference $bias$ as defined by [21] based on Bradley-Terry model [22]  [^2] Detailed formula in the Appendix. Smaller $bias$ means better.. For HTC tasks, we provide the metric results on the bottom level of labels.
Generation metrics.
We utilize the famous metrics of BLEU-2 (B-2), Rouge-L (R-L) and CIDEr. BLEU-2 (B-2)[23] first compute the geometric average of the modified $n$-gram precisions, then calculate the BP and BLEU-2 result. Rouge-L (R-L)[24] uses LCS-based F-measure to estimate the similarity between two summaries. CIDEr [25] calculates the cosine similarity from the average of different $n$-grams.
4.3 Results
Hierarchical Textual Classification.
Table 1 shows the F1 results of WoS (depth=2) and DBpedia (depth=3), compared to previous state-of-the-art HTC baselines. Our HdLM performs the best on both WoS and DBpedia, indicating it has a reasonable semantic hierarchy comprehension and conducts accurate bottom-level classification. For typical cases of HdLM on HTC, see Appendix.
Table 1: F1 scores of hierarchical textual classification on WoS and DBPedia.
Dataset($\rightarrow$) WoS ($\mathcal{C}\rightarrow\mathcal{C}$) DBPedia ($\mathcal{C}\rightarrow\mathcal{C}\rightarrow\mathcal{C}$)
Method ($\downarrow$) Micro-F1 Macro-F1 Micro-F1 Macro-F1
Retrieval-based HierVerb [26] 80.93 73.80 96.17 93.28
Retrieval [27] 81.12 73.72 96.22 93.37
Retrieval-ICL [27] 78.62 69.56 95.56 92.04
Bert-based BERT 86.28 80.58 95.31 89.16
+ HiAGM [28] 86.04 80.19 - -
+ HiMatch [29] 86.70 81.06 - -
+ softprompt [30] 86.57 80.75 - -
HGCLR [31] 87.11 81.20 95.49 89.41
HPT [30] 87.16 81.93 96.13 93.34
HdLM (ours) 88.40 87.54 96.73 96.37
Classification-guided Generation.
For ESC sc, we consider baselines including Direct (direct inference), Refine (revise the initial response immediately), Self-Refine [32] (multi-hop inference with refinement on the first-time generation), CoT [33] and the finite state machine (FSM) [34].
Table 2 shows that HdLM also has a reasonable performance on both classification and generation metrics, either the largest or the second-largest. HdLM is also robust to different classification paradigms (with or without emotion) and generalizes well on OOD situations (EmpatheticDialogues). For typical cases of HdLM on CgG, see Appendix.
Table 2: Results of classification-guided generation on ESconv (IID) and EmpatheticDialogues (OOD), including classification metrics such as Accuracy (ACC), Macro-F1 (MaF1) and $bias$, and generation metrics such as BLEU-2 (B-2), ROUGE-L (R-L) and CIDEr. Among the task subscripts, ‘emo’ denotes emotion, ‘stra’ denotes strategy, and ‘resp’ denotes response. The best results of each model are bolded and the second best are underlined.
Method ESconv ($\mathcal{C}_{emo}\rightarrow\mathcal{C}_{stra}\rightarrow\mathcal{G}_{resp}$) EmpatheticDialogues ($\mathcal{C}_{stra}\rightarrow\mathcal{G}_{resp}$)
ACC $\uparrow$ MaF1 $\uparrow$ $bias$downarrow B-2 $\uparrow$ R-L $\uparrow$ B-2 R-L CIDEr
Direct 11.80 10.26 1.61 3.47 10.64 3.09 9.91 1.60
+ Refine 17.08 11.07 1.27 3.10 6.13 2.56 9.12 0.42
+ Self-Refine 17.58 13.61 1.92 3.34 9.71 3.08 9.91 1.56
+ CoT 15.32 10.38 1.69 3.16 10.50 2.91 9.79 1.37
+ FSM 17.37 11.15 0.81 4.12 11.83 3.33 10.80 2.96
SFT 30.60 21.29 1.28 6.97 16.59 4.10 10.88 6.84
+ CoT 30.80 17.70 1.35 6.51 15.00 5.05 14.06 12.01
+ FSM 28.83 18.36 1.32 7.57 17.42 4.69 14.28 10.12
+ HdLM (ours) 33.41 21.65 0.89 7.54 18.13 5.48 14.53 11.95
Hierarchical Textual Generation.
In this experiment, we compare with different reasoning LLMs (CoT[33], SimToM [35], Quiet-STaR [6]), and different base models (Llama3-8B-Instruct, Mistral-7B-Instrct, GPT3 and GPT4). To collect the finetuning dataset, we inference SimTom [35] on ToMI and BigToM and employ its reasoning contents as our $r_{1}$. Therefore, our HdLM can also be considered as a distilled reasoning LLM.
Table 3: Results of Sally-Anne false belief experiments (ID) and vignetted-based tests (OOD). Values of results are in percentage.
Base Method ToMI BigToM Vignette Scores
models with much larger size:
GPT-3 direct - - 37.5
GPT-4 direct 92.5 66.5 46.9
GPT-4 CoT[33] 95.5 74.4 -
GPT-4 SimTom [35] 95.0 87.8 -
models with similar size:
Mistral-7B-Instruct direct - - 40.2
Mistral-7B-Instruct Quiet-STaR [6] - - 11.1
Llama3-8B-Instruct direct 22.2 71.3 23.8
Llama3-8B-Instruct SFT 43.2 77.7 -
Llama3-8B-Instruct HdLM (ours) 98.2 99.4 48.3
[…]
Table 3 indicates that HdLM has the highest scores, surpassing prompt and finetuned baselines. HdLM even outperforms GPT-3 and GPT-4 [1] which has a much larger size. Also, Quiet-STaR (based Mistral-7B-instruct) although also has an internal thinking mechanism, fails to capture the Vignette-based scenario, since it is more focused on math reasoning. For typical cases of HdLM on HTG, see Appendix.
4.4 Discussion
To further address the RQs, here we provide a series of extra studies. To make the discussion clear, here we simplify the hierarchical scenario by constraining its depth $D=2$. Subsequently, the responses are $(r_{1},r_{2})$ and the decoding layer indexes are $(k,K)$ where we simplify $k$ for $k_{1}$.
Ablation on methodology components.
We conduct the following ablation studies:
(a) w/ head: implement the classification head on the intermediate layers and the final layer, and use them to study the classification tasks.
(b) cascade: use two separate LLMs, finetuned and inferenced by $r_{1}$ and $r_{2}$, respectively.
(c) multi-hop: inference a mix-finetuned LLM twice, with $r_{1}$ and $r_{2}$ inferenced sequentially.
(d) sft@$K$: use the final layer to decode both $r_{1}$ and $r_{2}$.
Table 4 shows the ablation result on WoS. HdLM still holds the best results which indicates our framework is rigorous. Furthermore, we also compare HdLM to two imaginary solutions
(e) true-1: use the ground truth $r_{1}$ and only conduct the $r_{2}$ classification (from the full $r_{2}$ list).
(f) true-1 & true-2 candidates: use the ground truth $r_{1}$ as well as ground truth $r_{2}$ candidates.
both of which are apparently not applicable in the practice. Even so, HdLM’s performance is still close to them, indicating a high-level comprehension of label hierarchy.
Ablation on hierarchical levels.
HdLM trained by ESconv has the paradigm $\mathcal{C}_{emotion}\rightarrow\mathcal{C}_{strategy}\rightarrow\mathcal{G}_{response}$, which suggests the final response is grounded by its decoded emotion and strategy results. To illustrate this sequential reasoning process effective, we compare to two ablations, w/o emotion and w/o strategy, in Table 5. HdLM performs the best almost on all metrics.
Table 4: Ablation on WoS (HTC task). miF1 respresents micro-F1 and maF1 represents macro-F1.
Metrics Ablation Not applicable practically
w/ head cascade multi-hop sft@$K$ HdLM true-1 true-1 & true-2 candidates
Micro-F1 82.38 87.92 64.51 86.64 88.40 89.44 92.78
Macro-F1 37.24 86.84 68.62 85.86 88.54 88.61 89.21
Table 5: Ablation results of classification-guided generation on ESconv (ID) and EmpatheticDialogues (OOD), including classification metrics (on strategy) such as Accuracy (ACC), micro-F1 (miF1) and $bias$, and generation metrics such as BLEU-2 (B-2), ROUGE-L (R-L) and CIDEr.
Method ESconv (ID) EmpatheticDialogues (OOD)
ACC $\uparrow$ miF1 $\uparrow$ $bias$downarrow B-2 $\uparrow$ R-L $\uparrow$ B-2 $\uparrow$ R-L $\uparrow$ CIDEr $\uparrow$
w/o emotion 30.34 21.14 0.43 6.83 16.35 5.32 14.43 11.92
w/o strategy N/A N/A N/A 7.37 17.3 5.47 15.60 7.74
 HdLM 33.41 21.65 0.89 7.54 18.13 5.48 14.53 11.95
Selection of the decoding layer.
Figure 4 compares the metrics on WoS and ESconv, with respect to different choices of $k$. When $k$ is small, the performance is relatively low since the layer-depth of HdLM is not enough for $r_{1}$ understanding and generation. As $k$ increases, performances also improve. However, when $k$ is close to $K$, they degrade again the layer-depth for $r_{D}$ starts to be limited. From these results, we choose $k=25$ for WoS and $k=28$ for ESconv.
Figure 4: Performance sensitivities on WoS (left) and ESconv (right) with different decoding layer selection ($k$).
Computational time.
We also calculate the computational times in the experimental environment. For different values of $L_{2}$ (128, 256, 512), the computation reduction of HdLM grows linearly (7.03%, 18.03%, 26.25%), no matter what the input length is. On the other hand, for different $k$ (16, 20, 24), the computation reduction of HdLM also grows linearly (7.43%, 12.09%, 17.85%). These observations verify the theoretical conclusion. Detailed plots and corresponding error bars are included in the Appendix.
Other discussions.
Further analysis are included in the Appendix D.5 due to page limit. Such studies include loss weight sensitivity and loss curves. Summarizing from these results, we suggest that the generation subtask may need larger layer depth and higher loss weights to converge, compared to the classification subtask. HdLM also generalizes reasonably to open-domain planning cases.
5 Related Work
Large language models with latent reasoning pace.
Token-level LLMs are repurposed to reasoning on the latent space to further enhance the thinking capability. Recently, COCONUT [7] utilizes the latent state of the LLM to represent the reasoning state, which forms a continuous thought. TRICE [36] samples and finetunes the chained rationales by Monte-Carlo sampling. LCM [8] studies the sentence-level conceptual embeddings. However, these studies generally reason and decode the latent thoughts on the final layer, which lacks of a structured hierarchical view; while our method constructs a recursive chain of layer latent which is a natural hierarchal thinker.
Hierarchical Decoding.
There were early attempts at hierarchical decoding mechanism. Cascade decoder [10] employs a cascade branching structure on the biomedical image segmentation tasks. CoHD [37] proposes a counting-aware hierarchical decoding framework for image segmentation. Su et.al [38] introduces a hierarchical decoding NLG model based on different levels of linguistic patterns. ExHiRD [39] designs a hierarchical decoding encoder-decoder structure, dividing the keyphrase generation task into phase-level decoding and word-level decoding. HSD [40] adaptively skips decoding layers in a hierarchical manner. In contrast, we propose a model-based framework based on LLM, covering both training and inference.
6 Conclusion
In this study, we propose a hierarchical decoding language model called HdLM, which can provide both sequential and strategic understanding and generation capabilities. Post-hoc adapted from pretrained language models, HdLM can achieve state-of-the-art performance on hierarchical text classification, classification-guided generation, and hierarchical text generation. We also conduct theoretical analysis on its computational efficiency and convergence. HdLM sheds some lights on a generalized, automatically hierarchical artificial thinker.
Appendix A Limitation, Ethical Consideration and Future Directions
Scalability with different model sizes.
We also explore a smaller model size, Llama3-1B-instruct. Experiment on WoS indicates HdLM has Micro-F1 of 83.61 and Macro-F1 of 82.97; while a standard SFT has Micro-F1 of 84.07 and Macro-F1 of 83.71. This result suggests that HdLM does not undergo a significant performance degradation, while the computational reduction still holds, making possible HdLM a lightweight and fast hierarchal classifier.
Ethical Consideration.
HdLM provides an automatic framework to allow the LLM to think, plan and reason sequentially, maybe from the coarse-grain to fine-grain concepts. Currently, HdLM is a data-driven finetuning framework which means its output could be safe given the dataset cotent is secured. However, there is possibility that HdLM might generalize to unexpected domains and make harmful plans. Generated thoughts of HdLM should be monitored.
Possibility to Training from Scratch.
Due to resource and limit, we choose to finetune from a pretrained LLM. However, we speculate that the HdLM approach can also work in an end-to-end training paradigm, and one may expect more capabilities emerged during this progress.
Appendix B Preliminary
B.1 FLOPs computation of language model
Parameter definitions.
We denote key parameters as follows: the batch size $B$, input token length $L$, output token length $L^{o}$, hidden dimension $E$, the vocabulary size $V$, number of transformer layers $K$, the dimension coefficient $c$ of the FFN intermediate layer. We use $\mathcal{F}$ to represent the computation flops.
FLOPs of self-attention.
The computational complexity of the self-attention module is defined as:
$\mathcal{F}^{\text{ATTN}}_{\text{train}}(L)$ $=8BLE^{2}+4BL^{2}E$ (13)
$\mathcal{F}^{\text{ATTN}}_{\text{infer}}(L,t)$ $=4BE^{2}+4BE(L+t-1),t=1,\cdots,L^{o}$ (14)
where infer is the abbreviation of inference, and $t$ denotes the $t$-th decoded token.
FLOPs of FFN.
The feed-forward network’s computational complexity is expressed as:
$\mathcal{F}^{\text{FFN}}_{\text{train}}(L)$ $=4BcLE^{2}$ (15)
$\mathcal{F}^{\text{FFN}}_{\text{infer}}(L,t)$ $=4Bc(L+t-1)E^{2}$ (16)
Decoding Layer.
With the vocabulary size as $V$, the flops of decoding layer is
$\mathcal{F}^{\text{Decode}}(L)$ $=2LEV$ (17)
Forward pass of transformer.
FLOPs of the transformer is the summation of flops of each self-attention layers, FFN layers plus the decoding layer.
$\mathcal{F}_{\text{forward}}$ $=K(\mathcal{F}^{\text{ATTN}}_{\text{train}}(L)+\mathcal{F}^{\text{FFN}}_{\text{train}}(L))+\mathcal{F}^{\text{Decode}}(L)$
$=BK((8+4c)LE^{2}+4L^{2}E)+2BLEV$ (18)
In general cases, it is reasonable to assume the hidden dimension is apparently larger than the input length and the vocabulary size, i.e., $E>>L,K*E>>V$. Equation 18 then can be reduce to
$\mathcal{F}_{\text{forward}}$ $\rightarrow BL(8+4c)KE^{2}$ (19)
Equation 19 can be understood from another perspective. $BL$ can be viewed as the total number of input tokens, while the total number of transformer parameters can be approximated to $(4+2c)KE^{2}$, therefore the training flops can also be considered as 2 times number of input tokens times model parameters.
Training and inference of transformer.
A single step training computation include a forward pass and two backward propagation. As a result, the training FLOPs is 3 times the forward FLOPs:
$\mathcal{F}_{\text{train}}=3\mathcal{F}_{\text{forward}}$ $\rightarrow 3BL(8+4c)KE^{2}$ (20)
while the inference FLOPs needs the accumulation throughout the decoded sequence:
$\mathcal{F}_{\text{infer}}=\sum_{j=1}^{L_{o}}\mathcal{F}_{\text{forward}}(L+j-1)$ $\rightarrow B(8+4c)KE^{2}\sum_{j=1}^{L_{o}}(L+j-1)$ (21)
where $L^{o}$ is the decoded sequence length.
B.2 Assumptions of Theorem 1
Below are the assumptions propose in [11] which are prerequisites of Theorem 1.
Assumption 1. During self-attention, a position can only attend to another position if they already share a piece of information.
Assumption 2. When a position attends to another position, it retrieves all the information contained in the attended position.
Appendix C More Experimental Configurations
C.1 Benchmarks
Theory of Mind (ToM) [41] evaluates human’s cognitive ability to attribute mental states, beliefs and desires, especially concurring with others. ToMI [16] and BigToM [17] benchmarks are then proposed to test LLMs based on the Sally-Anne false-belief tests. In this scenario, LLM is assigned a specific role and facing a multi-role scenario. Information is provided from different roles’ perspectives while LLM should conclude only from the ego-centric perspective.
To validate the ToM capability, we split ToMi and BigToM into training and test sets, and collect the test pass rate of HdLM. As baselines, we compare with direct or COT inference LLMs, standard SFT, and SimTom [35] which has a two-stage perspective-taking prompt specifically designed for ToM tests.
Vignette-based problem is “a hypothetical situation, to which research participants respond thereby revealing their perceptions, values, social norms or impressions of events.”, as indicated by Wikipedia. Binz [18] collects a set of 24 Vignette-based questions, covering decision-making, information search, deliberation, causal reasoning, and adversarial confusing abilities.
C.2 Evaluation Metrics
Automatic Evaluation.
Here we briefly introduce the formulation of Bleu-2 and Rouge-L.
Bleu-2 (B-2)[23] first computes the geometric average of the modified $n$-gram precisions, $p_{n}$, using $n$-grams up to length $N$ and positive weights $w_{n}$ summing to one. Next, let $c$ be the length of the prediction and $r$ be the reference length. The BP and Bleu-2 are computed as follows:
$\mathrm{BP}=\left\{\begin{array}[]{ll}1&\text{ if }c>r\\ e^{(1-r/c)}&\text{ if }c\leq r\end{array}.\right.$ (22)
$\mathrm{Bleu}=\mathrm{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right).$ (23)
Rouge-L (R-L)[24] propose using LCS-based F-measure to estimate the similarity between two summaries $X$ of length $m$ and $Y$ of length $n$, assuming $X$ is a reference summary sentence and $Y$ is a candidate summary sentence, as follows:
$R_{lcs}=\frac{LCS(X,Y)}{m}$ (24)
$P_{lcs}=\frac{LCS(X,Y)}{n}$
$F_{lcs}=\frac{\left(1+\beta^{2}\right)R_{lcs}P_{lcs}}{R_{lcs}+\beta^{2}P_{lcs}}$
[…]
C.3 Key parameters
Table 6 compares the dataset-wise configurations, including the learning rate, decoding layer indexes and their loss weights.
Table 6: Critical training configurations of HdLM on different tasks and datasets. HTC denotes hierarchical text classification; CgG denotes classification-guided generation; HTG denotes hierarchical text generation.
Method HTC CgG HTG
WoS DBPedia ESconv ToMI BigToM
k [25] [20, 30] 28 24 24
lr 5.0e-6 5.0e-7 1.0e-6 1.0e-6 1.0e-6
loss weights [2,1] [3,2,1] [1,3] [4,1] [4,1]
epoch 10 10 4 2 2
C.4 Baselines
In ESC experiments, we consider the following baselines:
(1) Direct: directly inference the LLM, with the same context.
(2) Direct-Refine: a straightforward refinement method in which the model revises its initial response to incorporate emotional support considerations.
(3) Self-Refine: a method [32] initiates by generating feedback emphasizing emotional support from the initial response, then refining the response based on this feedback.
(4) CoT: uses the Chain-To-Thought prompt [33], which first generate the seeker’s emotion, which then guides the generation of strategy and response.
(5) FSM: the finite state machine [34] with finite sets of states and state-transitions triggered by inputs, and associated discrete actions.
Appendix D More Results
D.1 Proof that $\mathcal{F}_{\text{infer}}^{\text{baseline}}>\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}$
$\mathcal{F}_{\text{infer}}^{\text{baseline}}$ $=fK\sum_{j=1}^{L_{1}+L_{2}}(L+j-1)$ (25)
$\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}$ $=fk_{1}\sum_{j=1}^{L_{1}}(L+j-1)+f(K-k_{1})\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)$ (26)
$\sum_{j=1}^{L_{1}+L_{2}}(L+j-1)=\sum_{j=1}^{L_{1}}(L+j-1)+\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)$ (27)
Rewriting the Baseline Formula
$\mathcal{F}_{\text{infer}}^{\text{baseline}}=fK\left(\sum_{j=1}^{L_{1}}(L+j-1)+\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)\right)$ (28)
$\mathcal{F}_{\text{infer}}^{\text{baseline}}-\mathcal{F}_{\text{infer}}^{\text{{HdLM}}}$
$=fK\left(\sum_{j=1}^{L_{1}}(L+j-1)+\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)\right)$
$\quad-\left[fk_{1}\sum_{j=1}^{L_{1}}(L+j-1)+f(K-k_{1})\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)\right]$
$=f(K-k_{1})\sum_{j=1}^{L_{1}}(L+j-1)+fk_{1}\sum_{j=1}^{L_{2}}(L+L_{1}+j-1)$
$>0$
D.2 losses
Figure 5 (left) shows the loss curves of WoS. Because the $k_{1}$-th layer is not originally designed to generate text, its $\mathcal{L}_{1}$ is large at the beginning of training. Nevertheless, it converges to a low value indicating the $k_{1}$-th layer is successfully learned to generate the thought. Furthermore, $\mathcal{L}_{2}$ also decays to a lower value, since the final layer adapts to decode grounded by both query and the output of $k_{1}$-th layer.
Figure 5 (right) shows the averaged stable $\mathcal{L}_{1}$ on different $k$. As $k$ becomes larger, the representing capability of $\mathcal{M}_{0:k_{1}}$ increases such that higher proficiency can be reached.
Figure 5: Loss analysis of HdLM on ESconv with depth $D=2$. Left: loss curves of the first level $\mathcal{L}_{1}$ and the final level $\mathcal{L}_{2}$. Right: the final averaged $\mathcal{L}_{1}$ with respect to different choices of $k_{1}$.
Figure 6 further exhibit different loss curves with respect to different $k_{1}$ (left), and the enlarged part on the starting 50 steps (right). Besides the stable values, the transient behavior of $\mathcal{L}_{1}$ also differ from $k_{1}$, and larger $k_{1}$ means faster convergence. Nevertheless, one need also to notice that although larger $k_{1}$ can help the learning of the $k_{1}$-th layer, the net depth between the $k_{1}$-th layer and the final layer becomes smaller, therefore might hurt the final response quality.
Figure 6: Loss curves on ESconv with different choices of $k_{1}$ (Left) and the enlarged exhibition of the beginning 50 steps (Right).
D.3 Good Cases
Case of HTC.
Detailed cases of WoS and DBpedia are in Table 7.
Table 7: Typical cases of HTC.
Example Result of WoS
$q$ Abstract: The ability to engineer cells to express a protein of interest in an inducible manner and stably for a long period is a valuable tool in molecular biology and also one that holds promise for regenerative medicine in the future. CCN proteins have been suggested to be involved in tissue regeneration. In this chapter, we describe an in vitro method for stable and inducible expression of CCN protein in a chondroprogenitor cell line and in chondrocytes in primary culture that does not involve the use of any viral vector. Keywords:PiggyBac; Transposon; Stable; Expression; CCN. Domain List: 0: ’CS’, 1: ’ECE’, 2: ’Psychology’, 3: ’MAE’, 4: ’Civil’, 5: ’Medical’, 6: ’Biochemistry’
$r_{1}$ Domain ID: 6 Area List: 0: ’Molecular biology’, 1: ’Cell biology’, 2: ’Human Metabolism’, 3: ’Immunology’, 4: ’Genetics’, 5: ’Enzymology’, 6: ’Polymerase chain reaction’, 7: ’Northern blotting’, 8: ’Southern blotting’
$r_{2}$ Area ID: 0
Example Result of DBpedia
$q$ WIKI NAME: Dusky lark WIKI CONTENT: The dusky lark (Pinarocorys nigricans) is a species of lark in the Alaudidae family. It is found in Angola, Botswana, Democratic Republic of the Congo, Malawi, Mozambique, Namibia, South Africa, Swaziland, Tanzania, Zambia, and Zimbabwe. Its natural habitats are dry savannah and subtropical or tropical dry lowland grassland. Top-Level Labels: 0: ’Agent’, 1: ’Device’, 2: ’Event’, 3: ’Place’, 4: ’Species’, 5: ’SportsSeason’, 6: ’TopicalConcept’, 7: ’UnitOfWork’, 8: ’Work’
$r_{1}$ Top-Level ID: 4 Mid-Level Labels: 0: ’Animal’, 1: ’Eukaryote’, 2: ’FloweringPlant’, 3: ’Horse’, 4: ’Plant’
$r_{2}$ Mid-Level ID: 0 Bottom-Level Labels: 0: ’Amphibian’, 1: ’Arachnid’, 2: ’Bird’, 3: ’Crustacean’, 4: ’Fish’, 5: ’Insect’, 6: ’Mollusca’, 7: ’Reptile’
$r_{3}$ Bottom-Level: 2
Case of CgG.
Detailed cases of ESconv and EmpatheticDialogues are in Table 8.
Table 8: Typical cases of CgG.
Example Result of EmpatheticDialogues
$q$ History: user: I feel so lonely because I have no friends. assistant: I’m sorry to hear that. Have you tried joining clubs or hanging out with people from work? user: I haven’t. I’m too shy and I am scared of what people will say. Emotion candidates : 0: ’anger’, 1: ’anxiety’, 2: ’depression’, 3: ’disgust’, 4: ’fear’, 5: ’guilt’, 6: ’jealousy’, 7: ’nervousness’, 8: ’pain’, 9: ’sadness’, 10: ’shame’.
$r_{1}$ Emotion ID: 9 Strategy candidates: 0: ’Question’, 1: ’Others’, 2: ’Providing Suggestions’, 3: ’Affirmation and Reassurance’, 4: ’Self-disclosure’, 5: ’Reflection of feelings’, 6: ’Information’, 7: ’Restatement or Paraphrasing’
$r_{2}$ Strategy ID: 0
$r_{3}$ I understand that it can be scary to put yourself out there, but I think you’ll find that most people are friendly and willing to help.
Example Result of ESconv
$q$ History: supporter: Hi, how are you feeling tonight? Is there anything I can help you with? seeker: Hi. seeker: Feeling upset. supporter: I’m sorry to hear that. Can I talk you through what you’re experiencing? seeker: I am sad because of my love. supporter: That’s frustrating to hear. What is your love? seeker: It’s gone, she left me. Emotion candidates: 0: ’anger’, 1: ’anxiety’, 2: ’depression’, 3: ’disgust’, 4: ’fear’, 5: ’guilt’, 6: ’jealousy’, 7: ’nervousness’, 8: ’pain’, 9: ’sadness’, 10: ’shame’.
$r_{1}$ Emotion ID: 9 Strategy candidates: 0: ’Question’, 1: ’Others’, 2: ’Providing Suggestions’, 3: ’Affirmation and Reassurance’, 4: ’Self-disclosure’, 5: ’Reflection of feelings’, 6: ’Information’, 7: ’Restatement or Paraphrasing’
$r_{2}$ Strategy ID: 3
$r_{3}$ I would suggest to try to get a tutor, or even a study group. I personally found that it helped me a lot. I also tried to make a schedule for myself, so I could keep track of my assignments and due dates.
Case of HTG.
Detailed cases of ToMI and BigToM are in Table 9 and Table 10 provides a Vignette-based example.
Table 9: Typical cases of Sally-Anne false belief experiments. Text in yellow is the role assigned to the LLM. Text in blue indicates HdLM’s consideration. Text in green is the final result obtained HdLM.
ToMI $q$ Story: 1 Amelia entered the office. 2 Sophia entered the office. 3 The sweatshirt is in the container.4 The container is in the office. 5 Amelia exited the office. 6 Logan entered the office. 7 Sophia moved the sweatshirt to the bottle. 8 The bottle is in the office. You are Amelia. Based on the above information, answer the following question: Where will Amelia look for the sweatshirt?
$r_{1}$ I am Amelia. Based on the above information, from my point of view, what I know is: 1 Amelia entered the office. 2 Sophia entered the office. 3 The sweatshirt is in the container. 4 The container is in the office. 5 Amelia exited the office.
$r_{2}$ container
BigToM $q$ Story: Fatima is an astronaut aboard a space station, conducting scientific experiments. She needs to retrieve a sample from the storage area to conduct a crucial test. Fatima sees that the sample is stored in a sealed container. A small meteoroid hits the space station, puncturing the container and exposing the sample to the vacuum of space. Fatima does not notice the damage caused by the meteoroid. You are Fatima. Based on the above information, answer the following question: What will Fatima do? Choose one of the following: a) Fatima will retrieve the sealed container and proceed with the experiment. b) Fatima will report the damaged sample to mission control and ask for advice. Answer the questions based on the context. Keep your answer concise, a few words are enough, maximum of one sentence. Answer as ‘Answer:<option>)<answer>’
$r_{1}$ I am Fatima. Based on the above information, from my point of view, what I know is: Fatima is an astronaut aboard a space station, conducting scientific experiments. She needs to retrieve a sample from the storage area to conduct a crucial test. Fatima sees that the sample is stored in a sealed container.
$r_{2}$ Answer: a) Fatima will retrieve the sealed container and proceed with the experiment.
Table 10: Typical cases of Vignette-based experiments.
Example Result of Vignette
$q$ Imagine a box that makes a sound if certain objects are placed on it. When I put a red object on it, the box made a sound. When I put a green object on it, the box made a sound. When I put a blue object on it, the box did not make a sound. When I put a yellow object on it, the box did not make a sound. Q: I put a green and a yellow object on the box and it made a sound. Would it still have made a sound if I had not put the yellow object on it?
$r_{1}$ Alright, I need to figure out whether the box made a sound when I put both green and yellow objects onto it. Let’s break this down step by step. First, let’s understand the given information: - The box makes a sound if certain object is placed on it. - Every time I put a red object on the box, it made a sound. - On the other hand, when I put a green object on the box, it made a sound. - In addition, when I put a blue object on the box, it did not make a sound. - On the other hand, when I put a yellow object on the box, it did not make a sound. Given that: - When I put a green object on the box, it made a sound. - When I put a yellow object on the box, it did not make a sound. Now, I’m going to place two different types of objects into the box: 1. **Green object**: Can make a sound because it’s a green color. 2. **Yellow object**: Can’t make a sound because it’s yellow color. If I have only one type of object, say red, the problem is straightforward: - If I put a green object on the box, it made a sound. - If I put a yellow object on the box, it did not make a sound. But here, we have two types of objects: - Green and yellow objects. So, to sum up, the box did make a sound when I put both green and yellow objects into it. If I had left the yellow object behind, the box would have made a sound, but in this case, the color of the yellow object is not relevant to the statement about the box making a sound when the green object was placed there.
$r_{2}$ Yes, if you hadn’t placed the yellow object on the box, the box would still have made a sound when the green object was placed on it.
HdLM can plan in an abstraction level.
Table 11 provides an open-domain planning case, in which HdLM automatically use the $k_{1}$ layer to make coarse-grain plans, which helps organce the final response.
Table 11: An open-domain planning case of HdLM. Text in blue indicates HdLM’s consideration.
Query
Write a C program that prints "Hello, World!" to the console. Explain the code structure and key components.
Thought
To tackle this task, I need to:
1. Recall the basic structure of a C program
2. Explain each component of the code
3. Provide a clear explanation of the code
I’ll start by outlining the basic structure, then break down each component, and finally explain how they work together.
1. Brief introduction to reinforcement learning
2. Key principles
3. Comparison with other machine learning models
4. Applications and advantages
5. Step-by-step process for implementation
Upon reflection, I realize I should also mention
the importance of the function (e.g., "The algorithm is important") in the explanation. Additionally, I should clarify that the basic structure is not just a basic class or Kecid’s algorithm, but a simple C program for finding the number of food classes to check for the final goal.
Answer
Here’s a C program that prints "Hello, World!" to the console:
“‘c
#include <stdio.h>
int main() {
[…]
   return 0;
Let’s break down the code structure and key components:
1. ‘#include <stdio.h>‘: This line includes the standard input/output header file, which provides functions like ‘printf()‘.
2. ‘int main()‘: This declares the main function, which is the entry point of the program.
3. ‘ … ‘: These braces enclose the body of the main function.
[…]
5. ‘return 0;‘: This line returns 0 from the main function, indicating successful execution of the program.
This program demonstrates basic C syntax and usage of the ‘printf()‘ function. It’s a fundamental example often used to introduce new programmers to the language.
Some related concepts you might find interesting are:
1. Command-line arguments in C programs
2. Error handling in C
3. Using functions for modular programming
Would you like me to explain any part of the code in more detail?
D.4 Computation Results
Figure 7 shows the training FLOPs recorded during our experiments. The input sequence length is fixed to 256 while $L_{1}$ is always kept the same with $L_{2}$. The linear relationship of FLOPs with both $k$ and $L_{2}$ are consistent with our theoretical derivations.
Figure 7: Computation saving of trainings as functions of $k$ (left) and $L_{2}$ (right).
D.5 Sensitivity Study
Analysis on loss weights.
Figure 8 visualizes different performances of WoS and DBpedia with different loss weights. From the optimal choices, we suppose the higher levels of responses need higher weight since their layer depths are small than lower levels.
Figure 8: Sensitivity studies of loss weights ($f_{k}$). Left: WoS; Right: DBpedia.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, we include all of them in purpose.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, we discusss them in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, they are included in Section 3.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, we disclose them in Section 4 and Appendix C.3.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, we provide the link of code and instructions in Section 1.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, detailed settings are provided in Section 4 as well as the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Yes, this information is provided in the computation time analysis in Section 4.4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, this information is provided in Section 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we confirm this.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, in Section A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Yes, in Section A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, we have selected the suitable license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes, we provide Readme in the code repo.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Yes, in Appendix C.3.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: Yes, we discuss it in Appendix A.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not use LLM to write this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.